# Variational Zero-shot Multispectral Pansharpening

### v1

@@ Abstract
Pansharpening aims to generate a high spatial resolution multispectral image (HRMS) by fusing a low spatial resolution multispectral image (LRMS) and a panchromatic image (PAN). The most challenging issue for this task is that only the to-be-fused LRMS and PAN are available, and the existing deep learning-based methods are unsuitable since they rely on many training pairs. Traditional variational optimization (VO) based methods are well-suited for addressing such a problem. They focus on carefully designing explicit fusion rules as well as regularizations for an optimization problem, which are based on the researcher’s discovery of the image relationships and image structures. Unlike previous VO-based methods, in this work, we explore such complex relationships by a parameterized term rather than a manually designed one. Specifically, we propose a zero-shot pansharpening method by introducing a neural network into the optimization objective. This network estimates a representation component of HRMS, which mainly describes the relationship between HRMS and PAN. In this way, the network achieves a similar goal to the so-called deep image prior because it implicitly regulates the relationship between the HRMS and PAN images through its inherent structure. We directly minimize this optimization objective via network parameters and the expected HRMS image through iterative updating. Extensive experiments on various benchmark datasets demonstrate that our proposed method can achieve better performance compared with other state-of-the-art methods. The codes are available at https://github.com/xyrui/PSDip.
@@ I Introduction
The multispectral pansharpening methods are mainly divided into four classes, i.e., the component substitution (CS) based methods, the multiresolution analysis (MRA) based methods, the variational optimization (VO) based methods, and the deep learning (DL) based methods. The CS-based methods and the MRA-based methods have an extensive research history. Both of them directly model the expressions of the HRMS and thus have very concise forms and remarkably low runtime. Specifically, the CS-based methods focus on extracting spatial information of PAN and adding it to the LRMS. The MRA-based methods apply multi-resolution decomposition to extract high-frequency information from PAN and then add it to the upsampled LRMS. The VO-based methods take multispectral pansharpening as an image inverse problem. The HRMS is considered as a variable that should be optimized through a formulated optimization problem. Previous VO-based methods focus mainly on two points. One is about characterizing the HRMS structure. The other is to describe the relationship between HRMS and PAN. The two points are then incorporated into the optimization problem via diverse loss terms or regularizations. DL-based methods take advantage of the data-fitting power of deep neural networks and directly learn the mapping from LRMS and PAN to the expected HRMS with the help of a large number of training pairs. However, DL-based methods often face challenges such as overfitting and lack of robustness when encountering new data. To address these issues, unsupervised DL-based methods eliminate the reliance on HRMS data. Instead, most existing unsupervised methods require a sufficient number of LRMS/PAN training pairs to learn the underlying relationships between HRMS, LRMS, and PAN of a specified scene [5, 6, 7, 8, 9, 10, 11]. The training pairs are comprised of part of this scene. The rest of the part is used to evaluate the trained network. When facing a new scene, the network should be retrained using LRMS/PAN pairs from this scene. Since no HRMS is available, existing unsupervised methods mainly focus on designing efficient loss functions to optimize the networks. More recently, research has been conducted on zero-shot pansharpening methods. These methods’ fundamental and challenging aspect lies in having only a single available pair of LRMS and PAN for both training and testing. In Fig. 1 we illustrate the difference between supervised, unsupervised and zero-shot methods from the perspective of training and testing dataset. Besides, a key difference between zero-shot and traditional methods is that the former incorporates additional parameters that need to be optimized, such as those from a neural network.
@@
Compared with VO-based methods, zero-shot methods leverage the powerful learning ability of deep neural networks. Existing zero-shot methods usually contain many relatively independent parts, each of which is used to learn one specific kind of image feature. In this work, we aim to simplify the zero-shot pansharpening learning process and thus propose a method that is based on a single optimization framework. The proposed method leverages the VO-based method and the so-called deep image prior (DIP). The VO-based method has the advantage that they can flexibly and carefully describe the relationship between HRMS, LRMS and PAN, and then design suitable optimization problems. By taking pansharpening as an image inverse problem, the LRMS is considered as a degraded version of the corresponding HRMS by first blurring and then downsampling. However, how to characterize the degradation process from HRMS to PAN remains a challenging problem. In early research works, PAN is simply seen as the result of the linear combination of the HRMS bands. The assumption soon did not meet the higher expectation of fusion performance. A series of methods with further exploration were then developed [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], e.g., the linear relationship is considered on the gradient field. Recently, inspired by CS- and MRA-based methods, HRMS has also been represented using the hadmard product of a coefficient tensor and a transformed PAN that has the same size as HRMS [22]. If the coefficient tensor is a completely free component, such representation, however, is noninformative, which is the reason that it does not provide effective image information about HRMS. Unlike previous works that would predetermine the coefficient tensor, we propose to place a regularization term about it in an optimization objective. Then, the coefficient tensor together with the expected HRMS are both treated as the variables to be optimized. Such formulation achieves two goals. First, it delivers effective information from PAN to HRMS because the coefficient tensor is constrained. Second, the structure of HRMS is also indirectly regularized via the above representation. Based on the recent discovery that the coefficient tensor contains image-looking structures [22], we consider the DIP-type regularization. Specifically, the coefficient tensor is predicted by a neural network whose structures can naturally deliver the image prior. In this way, the network parameters and the expected HRMS are optimized by minimizing the optimization objective. We simply iteratively update network parameters and the expected HRMS using gradient descent-based methods.
@@
  • A zero-shot multispectral pansharpening method is proposed, which does not require additional training data other than the observed LRMS and PAN. Specifically, we formulate an optimization objective in which the regularization of the coefficient is contained. The coefficient and the expected HRMS are taken as variables to be optimized within the one optimization problem.
  • A DIP-type regularization is considered for the coefficient tensor in the formulated optimization objective. Specifically, we use a neural network to predict the coefficient and the network structure itself could provide necessary prior to reconstruct the coefficient. In this way, the network parameters are optimized together with the expected HRMS.
  • We propose an easy-to-implement strategy to solve the formulated optimization problem. Specifically, the network parameters and the expected HRMS are iteratively updated about their corresponding subproblems by gradient descent methods. To stabilize the optimization process, the network is first initialized before the iterative updating.
@@ II-A Multispectral pansharpening
Applying deep neural networks to solve multispectral pansharpening has become a popular trend since the remarkable development of deep learning. Supervised DL-based pansharpening methods usually achieve superior performance due to the carefully designed network structure and the large amount of training data that contains ground-truth HRMS [40, 41, 42, 43, 44, 45, 46, 47, 48]. However, supervised DL-based methods may suffer performance reduction when tested on a different dataset. Besides, training on the reduced resolution data, which is most available, does not necessarily lead to the same good performance on full resolution data [49]. Thus, unsupervised DL-based methods have been investigated [5, 6, 7, 8, 9, 10, 11]. They are largely based on the network structure, the combination of inputs, and the design of loss functions. The training label requirement is released, but a large number of LRMS/PAN training pairs are still needed. Testing is usually performed on the same dataset. Most recently, zero-shot pansharpening methods have been proposed, in which no additional training data is required. The method in [50] contains three components, that is, RSP to pre-train a network through the supervised learning way, SDE to learn the spatial relationship between HRMS and PAN, and finally FUG to train the fusion network using the results of RSP and SDE. [51] propose to solve a variational optimization problem. Its objective contains the output of a neural network, which is an approximation to the HRMS. The networks are optimized by a series of loss functions independent of the variational optimization problem. Unlike [51], our method is built within one variational optimization problem, which is more concise and easier to implement.
@@ III Main Method
Figure 2: An overview of the proposed PSDip. We formulate an optimization problem $\min_{\mathcal{X},\theta}\leavevmode\nobreak\ \|\mathcal{Y}-(\mathcal{X}\otimes K)\downarrow_{r}\|_{F}^{2}+\lambda\|\mathcal{X}-f_{\theta}(\mathcal{X},P)\odot\hat{P}\|_{F}^{2},$ for multispectral pansharpening, where $f_{\theta}(\mathcal{X},P)$ means the coefficient tensor $\mathcal{G}$ of representation $\mathcal{X}=\mathcal{G}\odot\hat{P}$. The optimization objective is denoted as $\mathcal{L}(\mathcal{X},\theta)$. At first, we initializing $f_{\theta}$ so that $f_{\theta}$ predicts a rough $\mathcal{G}$. Then, the network parameters and the expected HRMS are formally optimized in the proposed optimization problem by iterative updating. When the optimization finishes after $T$ steps, we derive the expected HRMS $\mathcal{X}_{T}$ as well as the coefficient tensor $\mathcal{G}_{T}=f_{\theta_{T}}(\mathcal{X}_{T},P)$.
@@ III-A Overview to pansharpening
Multispectral pansharpening aims to fuse the observed LRMS and PAN into the expected HRMS. Let $\mathcal{X}\in\mathbb{R}^{H\times W\times S}$ denote the HRMS, where $H,W$ and $S$ represent the height, width and spectral dimension of HRMS, respectively. The corresponding LRMS refers to as $\mathcal{Y}\in\mathbb{R}^{h\times w\times S}$ and PAN is written as $P\in\mathbb{R}^{H\times W}$. The downscaling factor of the LRMS is written as $r$, i.e., $H/h=W/w=r$. For VO-based methods, the multispectral pansharpening problem can be modelled by the following framework:
@@
The most adopted relationship between HRMS and LRMS in VO-based pansharpening methods is
@@
Unlike $L_{y}$, the term $L_{p}$ is more difficult to design because it remains an open problem how the PAN degenerates from the HRMS. In the earliest works, the simple linear model $P=\mathcal{X}\times_{3}\bm{\mathrm{p}}$ was proposed, which means performing a linear combination of the spectral bands of $\mathcal{X}$ with the coefficients vector $\bm{\mathrm{p}}\in\mathbb{R}^{S\times 1}$. Then, more complex models to describe HRMS/PAN have been investigated to improve the guidance of PAN. For example, a few works seek back to representations proposed by CS- and MRA-based methods. Then, HRMS can be modeled as
@@
Eq. (4) has a very concise and flexible form, which is similar to Brovey transform [24]. Only two components are included in this representation. The extended PAN $\hat{P}\in\mathbb{R}^{H\times W\times S}$ is constructed from the PAN and has the same size, especially the spatial resolution, as the HRMS. For example, $\hat{P}$ could be histogram-matched by the PAN and LRMS. We see that in this representation, $\hat{P}$ mainly provides spatial information for HRMS. However, since PAN only has one channel, $\hat{P}$ does not preserve consistent spectral information with $\mathcal{X}$. That is why we need a coefficient tensor $\mathcal{G}$ to balance the approximation. Following Eq. (4), the term $L_{p}$ can be formulated as
@@
Let us take a closer look to Eq. (5). Suppose $\mathcal{G}$ is completely undetermined, $L_{p}$ obviously can not be used for optimization problem (1) because whatever $\mathcal{X}$ is, we can always set $\mathcal{G}=\mathcal{X}\oslash\hat{P}$ so that $L_{p}$ reaches its minimum value, i.e., zero. In other words, $\mathcal{X}=\mathcal{G}\odot\hat{P}$, if without additional constraints, actually does not reveal effective image information about $\mathcal{X}$. Previously, in both CS-based, MRA-based and relevant VO-based methods, $\mathcal{G}$ should be preset and various ingenious approaches have been proposed. When $\mathcal{G}$ is determined by a certain method, we see that $\mathcal{G}\odot\hat{P}$ is a direct approximation to $\mathcal{X}$ and $L_{p}$ directly constrains the value of $\mathcal{X}$.
@@ III-B Proposed model
The main difference of (6) is that a regularization term about $\mathcal{G}$ is added. As mentioned above, if the regularization is directly about $\mathcal{X}$, $L_{p}=\|\mathcal{X}-\mathcal{G}\odot\hat{P}\|_{F}^{2}$ can always reach zero no matter what $\mathcal{X}$ is. Then, $\hat{P}$, which contains information from PAN, does not effectively contribute to the formulated problem. However, the regularization on $\mathcal{G}$ makes the spatial fidelity term $L_{2}$ affect the problem solution. In this way, problem (6) is not highly ill-posed and the expected HRMS can be effectively guided by LRMS and PAN. Besides, since $\mathcal{G}\odot\hat{P}$ is an approximation to $\mathcal{X}$, regularization on $\mathcal{G}$ can be seen as indirectly constraining the structure of $\mathcal{X}$. To see this more clearly, we may write $R(\mathcal{G})\approx R(\mathcal{X}\oslash\hat{P})$.
@@
As revealed by [22], the coefficient tensor $\mathcal{G}$ contains image structures although it is not strictly an “image”. An example is shown in Fig. 3. We can see clear textures existed in $\mathcal{G}$. These textures contain not only spatial details which make $\mathcal{G}$ look like an image but also spectrum compensation information to match $\hat{P}$ with $\mathcal{X}$. The presence of these structures in $\mathcal{G}$ is mainly attributable to the high spatial similarity between HRMS and PAN. Specifically, since the spatial resolution of HRMS and $\hat{P}$ should be the same, the positions of the low-frequency area and the high-frequency area between HRMS and $\hat{P}$ are nearly the same. Thus, the element-division result of them, i.e., $\mathcal{G}$, would give rise to similar “low-frequency” and “high-frequency” structures in the same positions. Based on this observation, a direct choice for $R(\mathcal{G})$ is regularization used for images. In this work, we propose to use a neural network $f_{\theta}(\cdot)$ to estimate the coefficient tensor $\mathcal{G}$. Following the idea of DIP, the network structure itself can implicitly regularize its output with implicit image prior. Then, we derive the proposed method which has the following optimization objective
@@
where we have $\mathcal{G}=f_{\theta}(\mathcal{X},P)$. The regularization $R(\mathcal{G})$ is absorbed in the network $f_{\theta}(\cdot)$ and does not appear explicitly. For the general setting of DIP, $f_{\theta}$ would take random noise as input. In this way, the network learns to construct the target purely from “nothing”. Here in problem (7), the network $f_{\theta}$ is designed to take $\mathcal{X}$ and $P$ as inputs. This is for three reasons. First, the representation $\mathcal{X}=\mathcal{G}\odot\hat{P}$ shows an evident relationship between $\mathcal{X},\mathcal{G}$ and $P$. Thus, it is reasonable to consider constructing a mapping from $\mathcal{X}$ and $P$ to $\mathcal{G}$. Second, these two inputs could provide additional information for $f_{\theta}$ to construct $\mathcal{G}$ except for the network structure, which we empirically find to be very useful to get better estimation. Third, we can dynamically and gradually modify $\mathcal{G}$ by adjusting the input $\mathcal{X}$ except for only optimizing the network parameters. We call the proposed problem (7) PSDip.

### v2

@@ Abstract
Pansharpening aims to generate a high spatial resolution multispectral image (HRMS) by fusing a low spatial resolution multispectral image (LRMS) and a panchromatic image (PAN). The most challenging issue for this task is that only the to-be-fused LRMS and PAN are available, and the existing deep learning-based methods are unsuitable since they rely on many training pairs. Traditional variational optimization (VO) based methods are well-suited for addressing such a problem. They focus on carefully designing explicit fusion rules as well as regularizations for an optimization problem, which are based on the researcher’s discovery of the image relationships and image structures. Unlike previous VO-based methods, in this work, we explore such complex relationships by a parameterized term rather than a manually designed one. Specifically, we propose a zero-shot pansharpening method by introducing a neural network into the optimization objective. This network estimates a representation component of HRMS, which mainly describes the relationship between HRMS and PAN. In this way, the network achieves a similar goal to the so-called deep image prior because it implicitly regulates the relationship between the HRMS and PAN images through its inherent structure. We directly minimize this optimization objective via network parameters and the expected HRMS image through alternating minimization. Extensive experiments on various benchmark datasets demonstrate that our proposed method can achieve better performance compared with other state-of-the-art methods. The codes are available at https://github.com/xyrui/PSDip.
@@ I Introduction
The multispectral pansharpening methods are mainly divided into four classes [5]. Among them, the component substitution (CS) based methods, the multiresolution analysis (MRA) based methods and the variational optimization (VO) based methods are also called traditional methods in this work. They do not contain learnable parameters and thus also do not require any training data. Deep learning (DL) based methods require large number of HRMS/LRMS/PAN training pairs to learn the neural network that generates HRMS. Unsupervised DL-based methods eliminate the reliance on HRMS. Instead, they use sufficient number of LRMS/PAN training pairs to learn the underlying relationships between HRMS, LRMS, and PAN of a specified scene [6, 7, 8, 9, 10, 11, 12]. The training pairs are comprised of part of this scene. The rest of the part is used to evaluate the trained network. When facing a new scene, the network should be retrained using LRMS/PAN pairs from this scene. More recently, zero-shot methods [13][14] have been researched. They also leverage neural network to learn the data relationship. However, they handle the most fundamental yet also challenging situation where only a single pair of LRMS and PAN is available for both training and testing. In Fig. 1 we illustrate the difference between supervised, unsupervised and zero-shot methods from the perspective of training and testing dataset.
This work handles the zero-shot situation that no other data except the observed LRMS and PAN is available. Following the VO-based works, we focus on modeling the HRMS/PAN relationship. Early VO-based works simply assume linear form in the spectral dimension. The assumption soon did not meet the higher expectation of fusion performance because 1) the linear assumption may not reflect the actual situation and 2) there still leaves much more image information that the linear assumption can not effectively provide. Thus, a series of methods with further exploration were then developed [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], e.g., the HRMS/PAN relationship is considered on the gradient field. Recently, inspired by CS- and MRA-based methods, HRMS has also been represented using the hadmard product of a coefficient tensor and a transformed PAN that has the same size as HRMS [25]. The biggest advantage of this representation is that, unlike linear form, it theoretically could have zero approximation error for any HRMS/PAN pairs. The problem is that if the coefficient tensor is a completely free component, such representation, however, is noninformative and does not provide effective image information about HRMS at all. In previous traditional works, the coefficient tensor is usually pre-estimated by different ingenious methods. The model performance largely relies on the setting of coefficient tensor. However, the pre-estimation process is usually independent from the optimization problem, which would limit the model performance.
@@
To address the above problem, we propose a new zero-shot multispectral pansharpening method which is based on a new proposed VO-based optimization framework. Specifically, we also represent HRMS by the product of a coefficient and a transformed PAN. But we place a regularization term about the coefficient in the general optimization objective for pansharpening. Then, the coefficient together with the expected HRMS are both treated as the variables to be optimized. Such formulation achieves two goals. First, it delivers effective information from PAN to HRMS because the coefficient is constrained. Second, the structure of HRMS is also indirectly regularized via the regularization on the coefficient. Actually, we proof that without changing the restored HRMS in the minimum point of the optimization problem, the regularization on HRMS can be completely absorbed into the regularization on the coefficient. Then, based on the recent discovery that the coefficient contains image-looking structures [25], we consider the DIP-type regularization [26]. Specifically, the coefficient tensor is predicted by a neural network whose structures can naturally deliver the image prior. In this way, the network parameters and the expected HRMS are optimized by minimizing the optimization objective. We simply iteratively update network parameters and the expected HRMS using gradient descent-based methods. Since the proposed method contains neural network that needs to be trained, it is a zero-shot method. Compared with existing zero-shot methods, the proposed method has more concise and compact form.
@@
  • A zero-shot multispectral pansharpening method is proposed, which does not require additional training data other than the observed LRMS and PAN. Specifically, we formulate a new optimization objective in which the regularization of the coefficient is contained. The coefficient and the expected HRMS are taken as variables to be optimized within the one optimization problem.
  • A DIP-type regularization is considered for the coefficient in the formulated optimization objective. Specifically, we use a neural network to predict the coefficient and the network structure itself could provide necessary prior to reconstruct the coefficient. Then, the network parameters are optimized together with the expected HRMS.
  • An easy-to-implement strategy, i.e., alternating minimization, is used to solve the formulated optimization problem. To stabilize the optimization process, the network is first initialized before the alternating minimization.
@@ II-A Multispectral pansharpening
Applying deep neural networks to solve multispectral pansharpening has become a popular trend since the remarkable development of deep learning. Supervised DL-based pansharpening methods usually achieve superior performance due to the carefully designed network structure and the large amount of training data that contains ground-truth HRMS [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. However, supervised DL-based methods may suffer performance reduction when tested on a different dataset. Besides, training on the reduced resolution data, which is most available, does not necessarily lead to the same good performance on full resolution data [57]. Thus, unsupervised DL-based methods have been investigated [6, 7, 8, 9, 10, 11, 12, 58, 59, 60]. They are largely based on the network structure, the combination of inputs, and the design of loss functions. The training label requirement is released, but a large number of LRMS/PAN training pairs are still needed. Testing is usually performed on the same dataset. Most recently, zero-shot pansharpening methods have been proposed, in which no additional training data is required. The method in [13] contains three components, that is, RSP to pre-train a network through the supervised learning way, SDE to learn the spatial relationship between HRMS and PAN, and finally FUG to train the fusion network using the results of RSP and SDE. [14] propose to solve a variational optimization problem. Its objective contains the output of a neural network, which is an approximation to the HRMS. The networks are optimized by a series of loss functions independent of the variational optimization problem. Unlike [14], our method is built within one variational optimization problem, which is more concise and easier to implement.
@@ III Main Method
Figure 2: An overview of the proposed PSDip. We formulate an optimization problem $\min_{\mathcal{X},\theta}\leavevmode\nobreak\ \|\mathcal{Y}-(\mathcal{X}\otimes K)\downarrow_{r}\|_{F}^{2}+\lambda\|\mathcal{X}-f_{\theta}(\mathcal{X},P)\odot\hat{P}\|_{F}^{2}$ for multispectral pansharpening. The optimization objective is denoted as $\mathcal{L}(\mathcal{X},\theta)$. The network $f_{\theta}(\mathcal{X},P)$ takes the HRMS $\mathcal{X}$ and PAN $P$ as inputs and outputs the coefficient tensor $\mathcal{G}$ which comes from the representation $\mathcal{X}=\mathcal{G}\odot\hat{P}$. We choose PanNet [45] as the backbone of $f_{\theta}$ (shown in the upper-left part). Before solving the optimization problem, we initialize $f_{\theta}$ by $\min_{\theta}\leavevmode\nobreak\ \|\hat{\mathcal{Y}}-f_{\theta}(\hat{\mathcal{Y}},P)\odot(\hat{P}\otimes K)\|$ so that $f_{\theta}$ predicts a rough $\mathcal{G}$ (shown in the upper-right part). Then, the network parameters and the expected HRMS are formally optimized in the proposed model by alternating minimization (shown in the lower part). Each subproblem is solved by gradient descent based methods. When the optimization finishes after $T$ steps, we derive the expected HRMS $\mathcal{X}_{T}$ as well as the coefficient tensor $\mathcal{G}_{T}=f_{\theta_{T}}(\mathcal{X}_{T},P)$.
@@ III-A Overview to pansharpening
Multispectral pansharpening aims to fuse the observed LRMS and PAN into the expected HRMS. Let $\mathcal{X}\in\mathbb{R}^{H\times W\times S}$ denote the HRMS, where $H,W$ and $S$ represent the height, width and spectral dimension of HRMS, respectively. The corresponding LRMS refers to as $\mathcal{Y}\in\mathbb{R}^{h\times w\times S}$ and PAN is written as $P\in\mathbb{R}^{H\times W}$. The downscaling factor of the LRMS is written as $r$, i.e., $H/h=W/w=r$. Taking multispectral pansharpening methods as an inverse problem [69], we can formulate the following model [70] for general VO-based methods [5]:
@@
The most adopted relationship between HRMS and LRMS in VO-based pansharpening methods is as follows [71]:
@@
Unlike $L_{y}$, the term $L_{p}$ is more difficult to design because it remains an open problem how the PAN degenerates from the HRMS. In the earliest works [38], the simple linear model $P=\mathcal{X}\times_{3}\bm{\mathrm{p}}$ was proposed and then used for many other works [39, 40, 41, 42, 43]. It means performing a linear combination of the spectral bands of $\mathcal{X}$ with the coefficients vector $\bm{\mathrm{p}}\in\mathbb{R}^{S\times 1}$. One of the biggest advantages of linear HRMS/PAN model is the computational convenience when solving the optimization problem because it only contains matrix multiplication operation. Besides, the spatial information can be well included since the operation is performed on the spectral dimension. However, the linear model has several drawbacks in improving the model performance. First, even if $P=\mathcal{X}\times_{3}\bm{\mathrm{p}}$ strictly holds, the solution space of this equation is quite large since $\mathrm{rank}(\bm{\mathrm{p}})=1$, meaning that there still exists much more image information that the linear model can not provide. Second, if the relationship $P=\mathcal{X}\times_{3}\bm{\mathrm{p}}$ does not strictly hold, such linear model would inevitably bring approximation errors.
In order to further improve the model performance., more complex models to describe HRMS/PAN have been investigated to improve the guidance of PAN [15, 16, 17, 21, 18, 19, 20, 22, 23, 24, 25]. For example, a few works seek back to representations proposed by CS- and MRA-based methods [23, 24, 25]. Among them, the most simple model represents HRMS by
@@
Eq. (4) has a very concise and flexible form, which is similar to Brovey transform [28]. Only two components are included in this representation. The extended PAN $\hat{P}\in\mathbb{R}^{H\times W\times S}$ is constructed from the PAN and has the same size, especially the spatial resolution, as the HRMS. For example, $\hat{P}$ could be histogram-matched by the PAN and LRMS. We see that in this representation, $\hat{P}$ mainly provides spatial information for HRMS. However, since PAN only has one channel, $\hat{P}$ does not preserve consistent spectral information with $\mathcal{X}$. That is why it needs a coefficient tensor $\mathcal{G}$ to balance the approximation. Theoretically, Eq. (4) could always hold no matter what real relationship between $\mathcal{X}$ and $P$ is. Following Eq. (4), the term $L_{p}$ can be formulated as
@@
Let us take a closer look to Eq. (5). Suppose $\mathcal{G}$ is completely undetermined, $L_{p}$ obviously can not be used for problem (1) because whatever $\mathcal{X}$ is, we can always set $\mathcal{G}=\mathcal{X}\oslash\hat{P}$ so that $L_{p}$ reaches its minimum value, i.e., zero. In other words, $\mathcal{X}=\mathcal{G}\odot\hat{P}$, if without additional constraints, actually does not reveal effective image information about $\mathcal{X}$. Previously, in both CS-based [28, 29, 30, 31, 32], MRA-based [34, 35, 36, 37] and relevant VO-based methods [23, 24, 25], $\mathcal{G}$ should be preset and various ingenious approaches have been proposed. When $\mathcal{G}$ is determined by a certain method, we see that $\mathcal{G}\odot\hat{P}$ is a direct approximation to $\mathcal{X}$ and $L_{p}$ directly constrains $\mathcal{X}$.
@@ III-B Proposed model
In the proposed model (6), a regularization term about $\mathcal{G}$ is added. As mentioned above, if the regularization is directly about $\mathcal{X}$, $L_{p}=\|\mathcal{X}-\mathcal{G}\odot\hat{P}\|_{F}^{2}$ can always reach zero no matter what $\mathcal{X}$ is. Then, $\hat{P}$, which contains information from PAN, does not effectively contribute to the formulated problem. However, the regularization on $\mathcal{G}$ makes the spatial fidelity term $L_{2}$ affect the problem solution. In this way, problem (6) is not highly ill-posed and the expected HRMS then tends to be more properly guided by both LRMS and PAN.
Besides, we simply remove $R(\mathcal{X})$ in model (6). Note that since $\mathcal{G}\odot\hat{P}$ is an approximation to $\mathcal{X}$, regularization on $\mathcal{G}$ can be seen as indirectly constraining the structure of $\mathcal{X}$. To see this more clearly, we may write $R(\mathcal{G})\approx R(\mathcal{X}\oslash\hat{P})$. Furthermore, we have the following theorem that implies that $R(\mathcal{X})$ can actually be “absorbed” into $R(\mathcal{G})$.
Theorem 1.
For any regularization $R_{g}(\mathcal{G})$ and $R_{x}(\mathcal{X})$, if $(\mathcal{X}^{*},\mathcal{G}_{1})$ is the minimum point of the following problem
$\min_{\mathcal{X},\mathcal{G}}\leavevmode\nobreak$ $\left\|\mathcal{Y}-(\mathcal{X}\otimes K)\downarrow_{r}\right\|_{F}^{2}+\lambda_{1}\left\|\mathcal{X}-\mathcal{G}\odot\hat{P}\right\|_{F}^{2}+$
$\lambda_{x}R_{x}(\mathcal{X})+\lambda_{g}R_{g}(\mathcal{G}),$ (7)
then, there exists at least one $R(\mathcal{G})$ and $\mathcal{G}_{2}$, such that $(\mathcal{X}^{*},\mathcal{G}_{2})$ is the minimum point of problem (6).
The proof is presented in Appendix A. With Theorem 1, our model (6) is somehow equivalent to model (1) with respect to $\mathcal{X}$. Comparing these two models, its seems like that $R_{x}(\mathcal{X})$ can be “absorbed” into $R_{g}(\mathcal{G})$ to form a new $R(\mathcal{G})$. Since the main focus of this work is about $\mathcal{G}$, we remove the regularization $R(\mathcal{X})$ to keep the model (6) as compact and concise as possible. Also we want to make our algorithm possibly easily reproducible. Considering this point, we prefer to keep the form of model (6) in this study.
@@
As revealed by [25], the coefficient tensor $\mathcal{G}$ contains image structures although it is not strictly an “image”. An example is shown in Fig. 3. We can see clear textures existed in $\mathcal{G}$. These textures contain not only spatial details which make $\mathcal{G}$ look like an image but also spectrum compensation information to match $\hat{P}$ with $\mathcal{X}$. The presence of these structures in $\mathcal{G}$ is mainly attributable to the high spatial similarity between HRMS and PAN. Specifically, since the spatial resolution of HRMS and $\hat{P}$ should be the same, the positions of the low-frequency area and the high-frequency area between HRMS and $\hat{P}$ are nearly the same. Thus, the element-division result of them, i.e., $\mathcal{G}$, would give rise to similar “low-frequency” and “high-frequency” structures in the same positions. Based on this observation, a direct choice for $R(\mathcal{G})$ is regularization used for images. In this work, we propose to use a neural network $f_{\theta}(\cdot)$ to estimate the coefficient tensor $\mathcal{G}$. Following the idea of DIP [26], the network structure itself can implicitly regularize its output with implicit image prior. Then, we derive the proposed method which has the following optimization objective
@@
where we have $\mathcal{G}=f_{\theta}(\mathcal{X},P)$. The regularization $R(\mathcal{G})$ is absorbed in the network $f_{\theta}(\cdot)$ and does not appear explicitly. For the general setting of DIP [26], $f_{\theta}$ would take random noise as input. In this way, the network learns to construct the target purely from “nothing”. Here in problem (8), the network $f_{\theta}$ is designed to take $\mathcal{X}$ and $P$ as inputs. This is for three reasons. First, the representation $\mathcal{X}=\mathcal{G}\odot\hat{P}$ shows an evident relationship between $\mathcal{X},\mathcal{G}$ and $P$. Thus, it is reasonable to consider constructing a mapping from $\mathcal{X}$ and $P$ to $\mathcal{G}$. Second, these two inputs could provide additional information for $f_{\theta}$ to construct $\mathcal{G}$ except for the network structure, which we empirically find to be very useful to get better estimation. Third, we can dynamically and gradually modify $\mathcal{G}$ by adjusting the input $\mathcal{X}$ except for only optimizing the network parameters. We call the proposed problem (8) PSDip.